How many liters each of a 40% acid solution and a 65% acid solution must be used to produce 60 liters of a 45% acid solution.

Answer by checkley79(3341)

.65X+.40(60-X)=.45*60
.65X+24-.40X=27
.25X=27-24
.25X=3
X=3/.25
X=12 LITERS OF 65% SOLUTION IS USED.
60-12=48 LITERS OF 40% SOLUTION IS USED.
PROOF:
.65*12+.40*48=.45*60
7.8+19.2=27
27=27